answersLogoWhite

0

ans- 4c3= 4!\3!*1!

=4*3*2*1\3*2*1

=4

User Avatar

Wiki User

∙ 17y ago

What else can I help you with?

Continue Learning about Statistics

A card is randomly drawn from an ordinary deck of playing cards what is the probability that the card is a king or a diamond?

You randomly select one card from a 52-card deck. Find the probability of selecting the king of diamonds or the jack of


You randomly select a month of the year What is the probability that you select a month with 28 or 30 days in it?

First we don't consider leap year There is 1 month with 28 days: february There are 4 months with 30 days: april, june, september, november So the probability to select a month with 28 or 30 days is (1+4)/12 = 5/12 If the year is a leap year then the probability is 4/12


A small business employs 6 women and 4 men. The manager randomly selects 3 of them to be on a team. What is the probability that all 3 are men?

1-30


What is the probability that 4 randomly selected people all have different birthdays?

Let us assume that there are exactly 365 days in a year and that birthdays are uniformly randomly distributed across those days. First, what is the probability that 2 randomly selected people have different birthdays? The second person's birthday can be any day except the first person's, so the probability is 364/365. What is the probability that 3 people will all have different birthdays? We already know that there is a 364/365 chance that the first two will have different birthdays. The third person must have a birthday that is different from the first two: the probability of this happening is 363/365. We need to multiply the probabilities since the events are independent; the answer for 3 people is thus 364/365 × 363/365. You should now be able to solve it for 4 people.


There are 5 yellow tiles 7 green tiles and 4 blue tiles in a bag What is the probability of randomly drawing 3 yellow tiles?

Try 5/16 x 4/15 x 3/14 = 1/56 = 0.010178571

Related Questions

A bag contains 3 red balls and 4 green balls randomly draw one ball then replace it and randomly draw a second what is the percentage?

Drawing one red ball out of a bag that contains 3 red balls and 4 green balls is 3 in 7. If we are to randomly select one ball out of the bag and replace it (I assume with the same ball) then there are still 3 red balls and 4 green balls. The odds are still 3 in 7. It doesn't matter how many times you draw. Every time you draw a ball you have a 3 in 7 chance of pulling out a red ball.


A card is randomly drawn from an ordinary deck of playing cards what is the probability that the card is a king or a diamond?

You randomly select one card from a 52-card deck. Find the probability of selecting the king of diamonds or the jack of


How do you boot people off of Halo Reach?

1. press start 2. go to players 3. select 4.kick


Four chips labeled 1 to 4 are placed in a hat Denise randomly selects three chips from the hat How many combinations of numbers are possible for Denise to select?

6


What is the shift pattern for a 13 speed?

Range Lo Selector back Low: 1, 1: 2, 2: 3, 3: 4, 4: 5 Range High 1: 6 select up 7, select down shift 2: 8 select up 9 select down shift 3: 10 select up 11, select down sift 4: 12 select up 13. Very Easy If you want Ill tell you 18 and 20 twin shift.


A professor must randomly select 4 students to participate in a mock debate. There are 18 students in his class. In how many different ways can these students be selected if the order of selection doe?

333


You randomly select a month of the year What is the probability that you select a month with 28 or 30 days in it?

First we don't consider leap year There is 1 month with 28 days: february There are 4 months with 30 days: april, june, september, november So the probability to select a month with 28 or 30 days is (1+4)/12 = 5/12 If the year is a leap year then the probability is 4/12


How do you change a text tone for vodafone 547?

1. Go to Profiles 2. Select the profile you require 3. Select customize 4. Select tone Settings 5 Select Message 6. select tone.


A small business employs 6 women and 4 men. The manager randomly selects 3 of them to be on a team. What is the probability that all 3 are women?

1-6


A small business employs 6 women and 4 men. The manager randomly selects 3 of them to be on a team. What is the probability that all 3 are men?

1-30


How many different ways are there to select a group of 4 out of 7 people for a seating arrangement in a row of length 4?

add ore people


What is the probability that 4 randomly selected people all have different birthdays?

Let us assume that there are exactly 365 days in a year and that birthdays are uniformly randomly distributed across those days. First, what is the probability that 2 randomly selected people have different birthdays? The second person's birthday can be any day except the first person's, so the probability is 364/365. What is the probability that 3 people will all have different birthdays? We already know that there is a 364/365 chance that the first two will have different birthdays. The third person must have a birthday that is different from the first two: the probability of this happening is 363/365. We need to multiply the probabilities since the events are independent; the answer for 3 people is thus 364/365 × 363/365. You should now be able to solve it for 4 people.

Trending Questions
What are the names of the queens in a deck of cards? A card is drawn at random from an ordinary pack of playing cards what is the probability that it is an ace? What is the average weight for an 6ft 16 year old male? Answers for For how many different counting numbers between 10 and 200 is the sum of the digits equal to 6 if zero is not a digit of any numbers? Possible to obtain a negative value for an F-ratio? What is your estimate for the total amount if time students in your school spent during this past semester studying for final exams? How effective are vasectomies after a long period of time? What is the probability of 6 any numbers in lottery of 40 numbers? What is the purposed normalization in database? What does the term initial risk level mean for hazards? Who invented the normal distribution? What does it mean when the distribution of data is skewed to the right? If you roll two fair dice what is the probability you will roll a double? How long after your period can you have unprotected sex? What is the sample space for picking a random card from a deck? What is the mass average of 10 aluminum cans? The distribution of the amount of change in UF student's pockets has an average of 2.02 dollars and a standard deviation of 3.00 dollars Suppose that a random sample of 45 UF students was taken and ea? What is the probability that there will be rain during at least one of three days? Is a graph an accurate way to represent data and when can you use it? The place where two or more railway lines meet?